# MaskCLIP++: A Mask-Based CLIP Fine-tuning Framework for Open-Vocabulary Image Segmentation

### v2

@@ Abstract
Open-vocabulary image segmentation has been advanced through the synergy between mask generators and vision-language models like Contrastive Language-Image Pre-training (CLIP). Previous approaches focus on generating masks while aligning mask features with text embeddings during training. In this paper, we observe that relying on generated low-quality masks can weaken the alignment of vision and language in regional representations. This motivates us to present a new fine-tuning framework, named MaskCLIP++, which uses ground-truth masks instead of generated masks to enhance the mask classification capability of CLIP. Due to the limited diversity of image segmentation datasets with mask annotations, we propose incorporating a consistency alignment constraint during fine-tuning, which alleviates categorical bias toward the fine-tuning dataset. After low-cost fine-tuning, combining with the mask generator in previous state-of-the-art mask-based open vocabulary segmentation methods, we achieve performance improvements of +1.7, +2.3, +2.1, +3.1, and +0.3 mIoU on the A-847, PC-459, A-150, PC-59, and PAS-20 datasets, respectively. Code is available at https://github.com/HVision-NKU/MaskCLIPpp.
@@ 1 Introduction
Image segmentation is one of the most extensively studied tasks in computer vision, which aims to partition an image into several regions where each region corresponds to the same object or shares consistent semantics. Traditional image segmentation models are often defined on a closed vocabulary. When new classes need to be segmented, it typically requires redefining the vocabulary set, annotating data, and retraining the model. However, the cost of annotating image segmentation on a large vocabulary is high. Therefore, the task of Open-Vocabulary Segmentation (OVS) has been proposed to perform image segmentation under any vocabulary set.
Large-scale image-text pre-training models, such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021), have demonstrated strong zero-shot recognition capabilities, sparking interest in transferring these abilities to image segmentation tasks. Several studies (Ding et al., 2023; Jiao et al., 2023; Yu et al., 2024; Jiao et al., 2024) have explored how to adapt mask generators (Cheng et al., 2021; 2022) and pre-trained CLIP to each other to achieve open-vocabulary segmentation at both the semantic and instance levels.
However, we argue that continuing to invest in improving the mask generation may not be cost-effective. For different CLIP architectures, even when the generated masks closely match the ground truth (GT) masks, the improvement in segmentation performance is limited to no more than 5% as shown in Figure 1(a). Conversely, for different mask generators, if CLIP can classify the generated masks as accurately as the GT class assignments, the segmentation performance improvement can exceed 40% as shown in Figure 1(b). Furthermore, Figure 1(a) shows that for the original CLIP, there remains a performance gap of 3.0%–4.7% between the generated masks and the GT masks, implying that continuing to use the generated masks during training may negatively impact the alignment quality between regional images and text. We confirm this inference in our experiments.
@@
The above analysis indicates that GT masks can replace mask generators during fine-tuning CLIP with image segmentation data, which not only eliminates unnecessary training costs but also improves the alignment quality between regional images and text. However, when attempting to use image segmentation data to fine-tune CLIP, a significant challenge is to avoid overfitting to the training set. Existing approaches either integrate a frozen CLIP during inference (Xu et al., 2022b; 2023a; Yu et al., 2024) or employ a frozen teacher model for distillation (Jiao et al., 2023; 2024). The former may limit the model’s performance improvement, while the latter often requires careful tuning of distillation strategies and hyperparameters. To address this, we analyze the sources of misalignment in previous methods and propose a consistency alignment constraint to ensure that alignment optimization preserves CLIP’s original embedding space. The outcome of our explorations is a new CLIP fine-tuning framework, coined as MaskCLIP++.
@@
Experiments show that MaskCLIP++ can significantly improve the mask classification capabilities of pretrained CLIP across multiple architectures, achieving 10.1%-16.8% mIoU improvement on ADE20K (Zhou et al., 2017). Additionally, MaskCLIP++ can be used with different mask generators to achieve open-vocabulary segmentation at the semantic- or instance-level. Compared to recent state-of-the-art methods, MaskCLIP++ offers lower training cost in terms of training time and memory usage, while achieving even better performance. We hope that the universality and efficiency of MaskCLIP++ could inspire further research in open-vocabulary image segmentation.
(a) Improvement potential of mask generation
(b) Improvement potential of mask classification
Figure 1: (a) Performance comparison of various original CLIP architectures using the same generated and ground truth masks. (b) Performance comparison of various mask generators with different backbones using CLIP based on ConvNeXt-L and ground truth classification. All models are trained on the COCO Panoptic dataset (Lin et al., 2014), with results reported on the ADE-150 (Zhou et al., 2017) dataset, measured by mean Intersection over Union (mIoU).
@@ 2 Related Work
Mask-based segmentation. Mask-based segmentation is a technique that views image segmentation as the process of mask generation and mask classification, which is also known as region-based segmentation. To address instance and panoptic segmentation, He et al. (2017); Wang et al. (2020; 2021); Zhang et al. (2021) proposed various mask generators. Recently, Cheng et al. (2021) unified the semantic, panoptic and instance segmentation into a single architecture. MaskFormer (Cheng et al., 2021) uses learnable queries to interact with image features, which are then decoded as masks and class scores. Mask2Former (Cheng et al., 2022) further accelerates the convergence of MaskFormer and reduces memory usage through multi-scale decoder design, using masks from shallow decoders as priors for deeper layers, and employing sampling-based optimization. Current mask-based OVS methods need almost to retrain an open-vocabulary mask generator based on Mask2Former (Xu et al., 2023d; a; Yu et al., 2024). In contrast, our method can directly use the Mask2Former trained on a closed vocabulary during inference, reducing training cost and increasing flexibility in usage.
Open-vocabulary segmentation. Depending on the data used, existing methods have developed into different paradigms. The first paradigm offers many heuristic solutions based on observations of CLIP without additional training (Zhou et al., 2022; Wang et al., 2024a; Li et al., 2023; Sun et al., 2024; Lan et al., 2024a; b). The second paradigm utilizes image-text pair data (Thomee et al., 2016; Changpinyo et al., 2021) to train OVS models under weak supervision (Xu et al., 2022a; Luo et al., 2023; Xu et al., 2023b; Zhang et al., 2023; Mukhoti et al., 2023; Cha et al., 2023; Liu et al., 2024a). Due to the lack of precise segmentation locations, these methods suffer from poor segmentation quality. The third paradigm enhances segmentation quality by incorporating image-mask data (Kirillov et al., 2023) and knowledge from SAM (Wang et al., 2024c; Yuan et al., 2024). However, the use of SAM not only increases inference cost but also relies on users or other detectors to provide high-quality prompts. The fourth paradigm seeks to achieve mutual adaptation between the segmentation model and CLIP through fully supervised training on image segmentation datasets containing image, mask, and category triplets, thereby producing an open-vocabulary segmentation model. Based on the type of the segmentation model, these methods can be categorized into mask-based approaches (Ding et al., 2022; Xu et al., 2022b; Ghiasi et al., 2022; Han et al., 2023; Liang et al., 2023; Qin et al., 2023; Jiao et al., 2023; Ding et al., 2023; Xu et al., 2023d; c; a; Yu et al., 2024; Liu et al., 2024b; Jiao et al., 2024) and pixel/patch-based ones (Li et al., 2022; Zhou et al., 2023; Li et al., 2024; Cho et al., 2024; Xie et al., 2024). Pixel-based OVS methods typically focus on semantic segmentation and cannot perform instance-level segmentation independently. Additionally, some unique approaches to achieving OVS involve image retrieval (Shin et al., 2022; Barsellotti et al., 2024), feature distillation (Chen et al., 2023; Wu et al., 2024) or model merging (Wang et al., 2024b). Some unified large models also incorporate the functionality of OVS (Zou et al., 2023; Shen et al., 2024). This paper focuses on efficiently transferring CLIP on limited segmentation datasets to improve its performance on semantic- and instance-level open-vocabulary segmentation tasks.
3 Methodology
3.1 Revisiting Image-Text Embeddings Alignment for OVS
We first revisit the roles of the visual encoder (CLIP-V) and text encoder (CLIP-T) in the pretrained CLIP model. Denote the image embeddings from CLIP-V as $E_{i}\in\mathbb{R}^{1\times D}$, where $D$ represents the dimension of embedding space. To create text embeddings, a vocabulary set of $K$ categories is transformed into the corresponding sentences using templates like “A photo of {}”, and then encoded by CLIP-T, yielding $E_{t}\in\mathbb{R}^{K\times D}$. The original CLIP model maps both images and text into an aligned embedding space. In this space, image-text pairs maintain semantic alignment, which we describe as “pre-alignment” and can often be measured using cosine similarity $\langle E_{i},E_{t}\rangle$. In open-vocabulary segmentation, preserving the pre-alignment property is crucial for adapting CLIP to zero-shot dense predictions, as it enhances the alignment of region-specific image and text embeddings.
Previous mask-based OVS methods obtain pre-aligned mask embeddings and text embeddings in two ways. As illustrated in Figure 2(a), the first way aims to train a mask generator that adapts to CLIP, with the generalization mainly derived from the original frozen parameters of CLIP (Xu et al., 2023c; Ding et al., 2023). However, attempts to align the embedding spaces of the mask generator and CLIP often fail (Xu et al., 2022b; Yu et al., 2024). As illustrated in Figure 2(b), the second way focuses on making CLIP sensitive to the generated masks during training to enhance performance and avoids overfitting through distillation (Jiao et al., 2023; 2024).
We identify two limitations of these two types of methods. First, they rely on noisy generated masks, which can degrade the quality of the region representations (as shown by the vertical comparisons in Figure 1(a)). This degradation can weaken the alignment between region-specific image and text embeddings. Second, there is a lack of an intuitive explanation and a straightforward solution for learning better alignment without severe overfitting. Moreover, freezing CLIP limits its transferability from global to local features, while distillation typically requires carefully designed strategies and incurs additional training cost.
To overcome these limitations, we propose a new training pipeline for OVS, named MaskCLIP++, which, to our knowledge, is the first method that attempts to leverage ground truth (GT) masks to obtain region features instead of relying on the mask generator during training. Furthermore, we propose the constraint named “consistency alignment” to alleviate overfitting when adding extra parameters to optimize alignment. We will give detailed descriptions in what follows.
@@ 3.1 Revisiting Image-Text Embeddings Alignment for OVS
Figure 2: Comparison of training pipeline between previous mask-based OVS methods. $\mathcal{L}$ denotes the loss function. (2(a)) adapts the mask generator to CLIP and avoids overfitting by freezing CLIP. (2(b)) adapts CLIP to the mask generator and avoids overfitting by distillation. “T” and “S” denote the teacher and student models, respectively. Our method (2(c)) abandons the mask generator and avoids overfitting by consistency alignment.
@@ 3.2 A New Fine-tuning Framework for CLIP
In this subsection, we detail how to use our framework to fine-tune CLIP and integrate it with an off-the-shelf mask generator. The overall framework is illustrated in Figure 3.
Extracting mask embeddings. Disregarding architectural details, CLIP-V can be considered as consisting of two parts: feature extraction by the encoder and feature aggregation by the neck. The mask is inserted after the last feature $F_{n}\in\mathbb{R}^{C\times H\times W}$ of the encoder is obtained, converting the original global aggregation into an aggregation within the mask’s effective region as follows:
$E_{m}=\operatorname{Proj}\sum_{i,j}\left({\phi(M)}_{i,j}\cdot{(F_{n})}_{:,i,j}\right),$ (1)
where $M$ is an object mask that has been downsampled to the same size as $F_{n}$, $E_{m}\in\mathbb{R}^{1\times D}$ is the embedding representations of this mask, $i,j$ are the spatial coordinates, and $\operatorname{Proj}$ represents the formulation of the neck that implements channel projection. The weighting function $\phi$ is related to the original CLIP pooling structure. Specifically, if the original CLIP uses global average pooling, $\phi$ acts as an L1 normalization function to achieve mask average pooling. If the original CLIP uses attention pooling, $\phi$ can be expressed as follows:
@@
$\phi(M)=\operatorname{Softmax}\left(\frac{q(T_{\text{cls}})k(T_{ij})}{\sqrt{d}}+\alpha M_{\text{bin}}\right),$ (2)
where $T_{\text{cls}}$ and $T_{ij}$ represent the global and local representations sent to the attention pooling, respectively. $q,k$ are the input projection of the attention. $d$ is the hidden size of attention. The spatially expanded and binarized version of $M$ is denoted as $M_{\text{bin}}$, where any elements less than $\max(M)/2$ are set to $\infty$- ∞.
Parameterized similarity modeling with consistency alignment. The pretrained CLIP mask embedding $E_{m}$ and the category text embedding $E_{t}$ are initially pre-aligned with the initial similarity map $S=\langle E_{m},E_{t}\rangle$ being relatively coarse. Optimizing this alignment with some supervision is needed to achieve a more accurate similarity map. We refer to this process as parameterized similarity modeling (PSM).
@@
However, when incorporating the mask generator into similarity modeling, the mask embeddings produced by the mask generator align almost exclusively with CLIP’s text embeddings on the training set, resulting in the misalignment illustrated in Figure 2(a). We first analyze the cause of this overfitting to develop our similarity modeling approach.
@@
To facilitate discussion, we simplify this approach as follows. Let $t\in\mathbb{R}^{D}$ denote a normalized text embedding from CLIP, and let $m_{1},m_{2}\in\mathbb{R}^{D}$ denote different mask embeddings. The initial similarity is defined as $s=m^{T}t$. Now, before computing the inner product across modalities, the mask embeddings are updated using a parameter $\Theta\in\mathbb{R}^{D\times D}$, resulting in a new similarity $r=(\Theta m)^{T}t$. Although $r$ can be optimized to be close to the target, an unintuitive phenomenon may arise: After mask embeddings are updated via gradient descent, the order of $r_{1}$ and $r_{2}$ is closer to the target, but the order of $s_{1}$ and $s_{2}$ may deviate further from the target. This phenomenon is illustrated by a toy example in Appendix Figure 5.
@@
Consequently, the original similarity relationships between CLIP’s embeddings are disrupted. Our experiments demonstrate that the additional parameter $\Theta$ fails to achieve CLIP’s level of generalization when fine-tuning on limited image segmentation data, resulting in overfitting. Therefore, when attempting to parameterize similarity modeling, it is essential to ensure that the modeled similarity maintains the same order as the original similarity, so that both can be optimized toward the same target. We refer to this constraint as “consistency alignment”.
To optimize under the consistency alignment constraint in most of the time, the simplest way to achieve this is to update the similarity map in a dimension that is irrelevant to any modality using linear layers. Specifically, for $Q$ masks and $K$ texts, the $Q\times K$ similarity matrix $S$ is first expanded to $Q\times K\times T$ using a linear layer of size $T$, and then it is reduced to $Q\times K$ using another linear layer of the same size.
When using the image patch embeddings $E_{p}$ to construct pseudo text embeddings (Jiao et al., 2024), a residual connection to the original text embeddings is necessary to ensure consistency alignment to some extent, as shown below:
$S=\langle E_{m},\operatorname{Norm}(E_{t}+P_{t})\rangle,\quad\text{where}~{}P_{t}=\operatorname{Attention}(q(E_{t}),k(E_{p}),v(E_{p}))$ (3)
Here, $\operatorname{Attention}$ denotes the scaled dot-product attention (Vaswani, 2017), $q,k,v$ denote the input projection functions for attention and $\operatorname{Norm}$ denotes the layer norm (Lei Ba et al., 2016). The image patch embeddings $E_{p}$ are equivalent to specialized mask embeddings extracted from unit masks on a per-patch basis.
Figure 3: The detailed framework of MaskCLIP++. The PSM represents the parameterized similarity modeling, which is constrained by the consistency alignment. The mask generator is only used during inference, and can be flexibly replaced.

### v3

# High-Quality Mask Tuning Matters for Open-Vocabulary Segmentation

@@ Abstract
Open-vocabulary image segmentation has been advanced through the synergy between mask generators and vision-language models like Contrastive Language-Image Pre-training (CLIP). Previous approaches focus on generating masks while aligning mask features with text embeddings during training. In this paper, we observe that relying on generated low-quality masks can weaken the alignment of vision and language in regional representations. This motivates us to present a new fine-tuning framework, named MaskCLIP++, which uses ground-truth masks instead of generated masks to enhance the mask classification capability of CLIP. Due to the limited diversity of image segmentation datasets with mask annotations, we propose incorporating a consistency alignment principle during fine-tuning, which alleviates categorical bias toward the fine-tuning dataset. After low-cost fine-tuning, MaskCLIP++ significantly improves the mask classification performance on multi-domain datasets. Combining with the mask generator in previous state-of-the-art mask-based open vocabulary segmentation methods, we achieve performance improvements of +1.7, +2.3, +2.1, +3.1, and +0.3 mIoU on the A-847, PC-459, A-150, PC-59, and PAS-20 datasets, respectively. Code is avaliable at https://github.com/HVision-NKU/MaskCLIPpp.
@@ 1 Introduction
(a) Performance: generated masks v.s. ground truth masks.
(b) Performance: CLIP classification v.s. ground truth classification.
Figure 1: Observations: 1(a) demonstrates the potential negative impact of low-quality generated masks on CLIP’s mask classification learning, while 1(b) showcases the untapped generalization potential of existing mask generators. Mask generators are trained on the COCO [36], with results reported on the ADE20K [80].
@@
Image segmentation is one of the most extensively studied tasks in computer vision, which aims to partition an image into several regions where each region corresponds to the same object or shares consistent semantics. Traditional image segmentation models are often defined on a closed vocabulary. When new classes need to be segmented, it typically requires redefining the vocabulary set, annotating data, and retraining the model. However, the cost of annotating image segmentation on a large vocabulary is high. Large-scale image-text pre-training models, such as CLIP [47] and ALIGN [25], have demonstrated strong zero-shot recognition capabilities, sparking interest in transferring these abilities to image segmentation tasks. Several studies [15, 26, 76, 27] have explored how to adapt mask generators [9, 10] and pre-trained CLIP to each other to achieve open-vocabulary segmentation at both the semantic and instance levels. However, our analysis based on two key observations reveals critical shortcomings in existing research that have been consistently overlooked.
First, we establish a baseline method enabling CLIP to perceive mask categories by referencing prior methodologies [73, 76]. To evaluate the impact of mask quality on inference performance, we conduct comparative experiments using generated versus ground-truth masks across various pre-trained CLIP models. As shown in Fig. 1(a), generated masks exhibit inferior performance compared to ground-truth masks due to their low segmentation accuracy, semantic ambiguity, and excessive redundancy. Our experiments further demonstrate that employing higher-quality masks during training consistently yields superior performance under identical inference protocols.
Second, to validate the strong generalization capability of existing mask generators, we evaluate multiple architectures pre-trained on the COCO dataset [36] and evaluate on the ADE20K dataset [80] with a different vocabulary set. Through IoU-based bipartite matching, we reassign “ground-truth categories” to predicted masks to determine whether these generators produce masks for unseen categories. As illustrated in Fig. 1(b), the performance could be substantially improved if mask classification were perfect. This discrepancy suggests that existing mask generators inherently produce numerous valid masks for unseen categories, but their potential is severely limited by erroneous mask classification.
@@
Based on these observations, we propose to eliminate the reliance on mask generators during training and directly fine-tune CLIP using high-quality image segmentation data. However, since image segmentation data is typically defined on closed vocabularies, mitigating the issue of CLIP overfitting the training data after fine-tuning poses an important challenge. Existing approaches either integrate a frozen CLIP during inference [72, 71, 76] or employ a frozen teacher model for distillation [26, 27]. The former may limit the model’s performance improvement, while the latter often requires careful tuning of distillation strategies and hyperparameters. To address this issue more effectively, we analyze the causes of modality misalignment in prior works and propose that the introduction of additional parameters should adhere to the consistent alignment principle, thereby preserving CLIP’s original embedding space throughout the alignment optimization process. The outcome of our explorations is a new CLIP fine-tuning framework, coined as MaskCLIP++.
@@
Experimental results demonstrate that MaskCLIP++ can significantly enhance CLIP’s masked classification capability across datasets spanning diverse domains. Additionally, MaskCLIP++ can be used with different mask generators to achieve open-vocabulary segmentation at the semantic- or instance-level. Compared to recent state-of-the-art methods, MaskCLIP++ offers lower training cost in terms of training time and memory usage, while achieving even better performance. We hope that the universality and efficiency of MaskCLIP++ could inspire further research in open-vocabulary segmentation.
@@ 2 Related Work
Mask-based segmentation. Mask-based segmentation is a technique that views image segmentation as the process of mask generation and mask classification, which is also known as region-based segmentation. To address instance and panoptic segmentation, [22, 62, 60, 79] proposed various mask generators. Recently, [9] unified the semantic, panoptic and instance segmentation into a single architecture. MaskFormer [9] uses learnable queries to interact with image features, which are then decoded as masks and class scores. Mask2Former [10] further accelerates the convergence of MaskFormer and reduces memory usage through multi-scale decoder design, using masks from shallow decoders as priors for deeper layers, and employing sampling-based optimization. Current mask-based OVS methods need almost to retrain an open-vocabulary mask generator based on Mask2Former [74, 71, 76]. In contrast, our method can directly use the Mask2Former trained on a closed vocabulary during inference, reducing training cost and increasing flexibility in usage.
@@
Figure 2: Comparison of training pipeline between previous mask-based OVS methods. $\mathcal{L}$ denotes the loss function. 2(a) adapts the mask generator to CLIP and avoids overfitting by freezing CLIP. 2(b) adapts CLIP to the mask generator and avoids overfitting by distillation. “T” and “S” denote the teacher and student models, respectively. Our method 2(c) abandons the mask generator and avoids overfitting by consistency alignment.
Open-vocabulary segmentation. Depending on the data used, existing methods have developed into different paradigms. The first paradigm offers many heuristic solutions based on observations of CLIP without additional training [81, 59, 33, 55, 29, 30]. The second paradigm utilizes image-text pair data [57, 7] to train OVS models under weak supervision [69, 40, 70, 78, 44, 6, 39]. Due to the lack of precise segmentation locations, these methods suffer from poor segmentation quality. The third paradigm enhances segmentation quality by incorporating image-mask data [28] and knowledge from SAM [63, 77]. However, the use of SAM not only increases inference cost but also relies on users or other detectors to provide high-quality prompts. The fourth paradigm seeks to achieve mutual adaptation between the segmentation model and CLIP through fully supervised training on image segmentation datasets containing image, mask, and category triplets, thereby producing an open-vocabulary segmentation model. Based on the type of the segmentation model, these methods can be categorized into mask-based approaches [14, 72, 19, 20, 35, 46, 26, 15, 74, 73, 71, 76, 38, 27] and pixel/patch-based ones [31, 82, 34, 11, 68]. Pixel-based OVS methods typically focus on semantic segmentation and cannot perform instance-level segmentation independently. Additionally, some unique approaches to achieving OVS involve image retrieval [52, 1], feature distillation [8, 65] or model merging [61]. Some unified large models also incorporate the functionality of OVS [83, 51]. This paper focuses on efficiently transferring CLIP on limited segmentation datasets to improve its performance on semantic- and instance-level open-vocabulary segmentation tasks.
3 Method
3.1 Revisiting Image-Text Embeddings Alignment for OVS
We first revisit the roles of the visual encoder (CLIP-V) and text encoder (CLIP-T) in the pretrained CLIP model. Denote the image embeddings from CLIP-V as $E_{i}\in\mathbb{R}^{1\times D}$, where $D$ represents the dimension of embedding space. To create text embeddings, a vocabulary set of $K$ categories is transformed into the corresponding sentences using templates like “A photo of {}”, and then encoded by CLIP-T, yielding $E_{t}\in\mathbb{R}^{K\times D}$. The original CLIP model maps both images and text into an aligned embedding space. In this space, image-text pairs maintain semantic alignment, which we describe as “pre-alignment” and can often be measured using cosine similarity $\langle E_{i},E_{t}\rangle$. In open-vocabulary segmentation, preserving the pre-alignment property is crucial for adapting CLIP to zero-shot mask classification, as it enhances the alignment of mask and text embeddings.
Previous mask-based OVS methods obtain pre-aligned mask embeddings and text embeddings in two ways. As illustrated in Fig. 2(a), the first way aims to train a mask generator that adapts to CLIP, with the generalization mainly derived from the original frozen parameters of CLIP [73, 15]. However, attempts to align the embedding spaces of the mask generator and CLIP often fail [72, 76]. As illustrated in Fig. 2(b), the second way make CLIP sensitive to the quality of generated masks during training to enhance performance and avoids overfitting through distillation [26, 27].
We identify two limitations of these two types of methods. First, they rely on noisy generated masks, which can degrade the quality of the region representations (as shown by the vertical comparisons in Fig. 1(b)). This degradation can weaken the alignment between region-specific image and text embeddings. Second, there is a lack of an intuitive explanation and a straightforward solution for learning better alignment without severe overfitting. Moreover, freezing CLIP limits its transferability from global to local features, while distillation typically requires carefully designed strategies and incurs additional training cost.
To overcome these limitations, we propose a new training pipeline for OVS, named MaskCLIP++, which, to our knowledge, is the first method that attempts to leverage ground truth (GT) masks to obtain region features instead of relying on the mask generator during training. Furthermore, we propose the principle named “consistency alignment” to alleviate overfitting when adding extra parameters to optimize alignment. We will give detailed descriptions in what follows.
Figure 3: Detailed framework of MaskCLIP++ for OVS tasks. The PSM represents the parameterized similarity modeling, which is designed under the principle of consistency alignment. The mask generator is only used during inference, and can be flexibly replaced.
@@ 3.2 A New Fine-tuning Framework for CLIP
In this subsection, we detail how to use our framework to fine-tune CLIP and integrate it with an off-the-shelf mask generator. The overall framework is illustrated in Fig. 3.
Extracting mask embeddings. Taking CLIP-V with an $L$-layer Transformers as an example, inspired by SAN [73], we posit that the layers from $1$ to $K$ predominantly perform feature extraction (denoted as the Extractor), while layers from $K+1$ to $L$ primarily conduct spatial fusion (referred to as the Fuser).
For the $N$ image tokens $F^{(l)}\in\mathbb{R}^{N\times C}$ that input to the $l$-th layer of the fusion module, we introduce $Q$ flattened image masks $M\in[0,1]^{Q\times N}$ as conditions to transform the fusion scope from global to local. As formally delineated in Eq. 1, we compute mask-conditioned modulation weights $\phi^{l}$ to perform weighted aggregation on $F^{(l)}$, followed by the layer-wise channel projection operator $\mathrm{Proj}^{(l)}$, thereby generating mask-specific tokens $E_{m}^{(l+1)}\in\mathbb{R}^{Q\times C}$.
$E_{m}^{(l+1)}=\operatorname{Proj}^{(l)}\left(\phi^{(l)}(M)\cdot F^{(l)}\right),~{}\text{where}~{}K<l<L.$ (1)
By leveraging the input projections $q$ and $k$ from the original attention module, we incorporate the mask into the weighting function through Eq. 2, thereby constraining the weight distribution within the mask’s active regions.
@@
$\phi^{(l)}(M)=\operatorname{softmax}\left(\frac{q(E_{m}^{(l)})k(F^{(l)})^{T}}{\sqrt{d}}+\alpha M_{\text{th}}\right),$ (2)
where $\sqrt{d}$ is the attention scaling factor, $\alpha\in(0,+\infty)$ is an learnable parameter, $M_{\text{th}}$ is the result of setting all values in $M$ that are less than $\max(M)/2$ to $\infty$- ∞.
When initializing $E_{m}^{(K+1)}$ with the CLS token in layer $K+1$, $E_{m}^{(L)}$ and $E_{t}$ exhibit pre-aligned properties. The first two rows of Tab. 1 quantitatively demonstrate the mask classification performance of pre-trained CLIP under this initialization paradigm.
Fig. 4 visualizes the normalized weight function $\phi$, which reveals the contributions of image tokens at different spatial positions to the generation of the current mask representation. After fine-tuning, more image tokens are incorporated into the construction of their corresponding mask representations.
Figure 4: Visualization of the weight function. The regions of interest are enclosed by yellow contours, and the deeper the red color at a position, the higher its importance to the region.
Parameterized similarity modeling with consistency alignment. Given the limited capacity of the original CLIP model in characterizing localized regions, a straightforward approach involves fine-tuning the original CLIP parameters alongside additional learnable parameters to reformulate the similarity metric between $E_{m}$ and $E_{t}$. We introduce a class of methods termed Parameterized similarity modeling (PSM) for this purpose.
@@
However, when incorporating the mask generator into similarity modeling, the mask embeddings produced by the mask generator align almost exclusively with CLIP’s text embeddings on the training set, resulting in the misalignment illustrated in Fig. 2(a). We first analyze the cause of this overfitting to develop our similarity modeling approach.
@@
To facilitate discussion, we simplify this approach as follows. Let $t\in\mathbb{R}^{D}$ denote a normalized text embedding from CLIP, and let $m_{1},m_{2}\in\mathbb{R}^{D}$ denote different mask embeddings. The initial similarity is defined as $s=m^{T}t$. Now, before computing the inner product across modalities, the mask embeddings are updated using a parameter $\Theta\in\mathbb{R}^{D\times D}$, resulting in a new similarity $r=(\Theta m)^{T}t$. Although $r$ can be optimized to be close to the target, an unintuitive phenomenon may arise: After mask embeddings are updated via gradient descent, the order of $r_{1}$ and $r_{2}$ is closer to the target, but the order of $s_{1}$ and $s_{2}$ may deviate further from the target. This phenomenon is illustrated by a toy example in Sec. A.2 in supplementary material.
@@
Consequently, the original similarity relationships between CLIP’s embeddings are disrupted. Our experiments demonstrate that the additional parameter $\Theta$ fails to achieve CLIP’s level of generalization when fine-tuning on limited image segmentation data, resulting in overfitting. Building upon the aforementioned analysis, we identify inconsistent alignment phenomena in PSM that prone to overfitting. This implies that preserving CLIP’s original similarity ordering to the greatest extent possible during optimization enables more effective utilization of additional parameters for alignment enhancement. We formalize this critical finding as the consistency alignment principle in PSM.
Specifically, for all mask embeddings $E_{m}\in\mathbb{R}^{Q\times C}$ extracted from an image and the text embeddings $E_{t}\in\mathbb{R}^{K\times C}$ of $K$ dataset categories, their similarity matrix $S=E_{m}E_{t}^{T}$ was projected through an over-parameterized linear projection along a newly created orthogonal dimension relative to both $Q$ and $K$ dimensions. Finally, cross-entropy loss was computed based on category annotations of each mask, guided by the optimized similarity map.